The heat capacity of aluminum is 0.900 J/g degrees Celsius . How much energy is needed to raise the temperature of a 8.50 x 10^2 g block of aluminum from 22.8 degrees Celsius to 94.6 degrees Celsius 


1
Expert's answer
2021-03-12T06:03:53-0500

E = m*c*(t2 - t1) = 8.5*10^2*0.9*(94.6 - 22.8) = 54927 J